Using the following data, which is the correct rate law of the sample reaction?
A + 5B + 6C → 3D + 3E
Experiment [A] (M) [B] (M) [C] (M)
Initial Rate
(M/s)
1 0.35 0.35 0.35 8x10-4
2 0.70 0.35 0.35 3.2x10-3
3 0.70 0.70 0.35 6.4x10-3
4 0.70 0.35 0.70 3.2x10-3
1
Expert's answer
2020-10-16T08:56:31-0400

Ans.

For this given reaction, A + 5B + 6C → 3D + 3E

The rate law may be R = k[A]a[B]b[C]c

R2 = 3.2 x 10-3 (M/S)  when [A] = 0.70 M, [B] = 0.35 M, [C] = 0.35 M

R3 = 6.4 x 10-3 (M/S) when [A] = 0.70 M, [B] = 0.70 M, [C] = 0.35 M

3.2 x 10-3  = k[0.70]a[0.35]b[0.35]c       (i)

6.4 x 10-3  = k[0.70]a[0.70]b[0.35]c   (ii)

Dividing equation ii by equation i


(0.70/0.35)b = 2 ; b = 1

R2 = 3.2 x 10-3 (M/S)  when [A] = 0.70 M, [B] = 0.35 M, [C] = 0.35 M

R4 = 3.2 x 10-3 (M/S) when [A] = 0.70 M, [B] = 0.35 M, [C] = 0.70 M

3.2 x 10-3  = k[0.70]a[0.35]b[0.35]c       (i)

6.4 x 10-3  = k[0.70]a[0.35]b[0.70]c   (iii)

Dividing equation iii by equation i

(0.70/0.35)c = 1 ; c = 0


R1 = 8 x 10-4 (M/S)  when [A] = 0.35 M, [B] = 0.35 M, [C] = 0.35 M

R2 = 3.2 x 10-3 (M/S)  when [A] = 0.70 M, [B] = 0.35 M, [C] = 0.35 M

3.2 x 10-3  = k[0.70]a[0.35]b[0.35]c       (i)

8 x 10-4  = k[0.35]a[0.35]b[0.35]c   (iv)

Dividing equation i by equation iv


(0.70/0.35)a = 4 ; a = 2

So the rate law for this reaction is R = k[A]2[B]1[C]0 (where k is rate constant) (Ans.)